OpenStudy (anonymous):

Find the equation of the tangent line to the curve y=(x+1)/(x-1) at the point (2,3)? CAn anyone help me please?

OpenStudy (anonymous):

take the derivative [(x+1)'(x-1) -(x+1)(x-1)' ]/(x-1)^2 [x-1 - x -1]/(x-1)^2 -2/(x-1)^2 so plug in your point y'(2) = -2/(2-1)^2 = -2/1^2 = -2 so thats your slope now use equation of a line y=mx+b 3 = -2(2) + b 3 = -4 + b so b = 7 thus the equation of the tangent line is y = -2x + 7